the total receipts from high school musical was $1274. Student tickets costs $2 each and adult tickets sold for $5 each. if they sold of 325 tickets that night, how many adult tickets were sold? many student tickets.
There were student tickets and adult tickets sold.

two cities are 16040 miles apart. a plane leaves one of them, traveling towards the other at an average speed of 410 miles per hour. At the same time a plane leaves the other, traveling towards the first, at an average speed of 410 miles per hour. How long will it take them to meet? the two meet in hours.

Answer by lwsshak3(11628)

The total receipts from high school musical was $1274. Student tickets costs $2 each and adult tickets sold for $5 each. if they sold of 325 tickets that night, how many adult tickets were sold? How many student tickets were sold?
..
let x=number of adult tickets sold
325-x=number of student tickets sold
5x+2(325-x)=1274
5x+650-2x=1274
3x=624
x=208
325-x=117
number of adult tickets sold=208
number of student tickets sold=117
..
Two cities are 16040 miles apart. a plane leaves one of them, traveling towards the other at an average speed of 410 miles per hour. At the same time a plane leaves the other, traveling towards the first, at an average speed of 410 miles per hour. How long will it take them to meet? the two meet in hours.
..
The two planes are traveling toward each other at a speed=410+410=820 mph
travel time=distance/speed=16040/820=19.56
How long will it take them to meet? 19.56 hrs